Integration by Partial Fractions


1.) ∫4dx/(x^4-1)


2.) ∫8dx/x(x^2+2)^2



1
Expert's answer
2021-05-07T09:12:31-0400

1) Solution:

"\frac{4}{x^4-1} = \frac{4}{(x-1)(x+1)(x^2+1)} = \frac{A}{x-1} + \frac{B}{x+1} + \frac{C}{x^2+1} = \newline
\frac{A(x+1)(x^2+1) + B(x-1)(x^2+1) + C(x-1)(x+1)}{(x-1)(x+1)(x^2+1)} = \newline
\frac{A(x^3+x^2+x+1) + B(x^3-x^2+x-1) + C(x^2-1)}{(x-1)(x+1)(x^2+1)} = \newline
\frac{x^3(A+B)+x^2(A-B+C)+x(A+B)+(A-B-C)}{(x-1)(x+1)(x^2+1)}"

from nominator

"A+B=0 \newline
A-B+C=0 \newline
A+B=0 \newline
A-B-C=4"

solution of this system is: A=1 B=-1 C=-2, so

"\frac{4}{x^4-1} = \frac{1}{x-1} - \frac{1}{x+1} - \frac{2}{x^2+1}"

and

"\int \frac{4}{x^4-1} dx = \int \frac{1}{x-1} dx - \int \frac{1}{x+1} dx - 2 \int \frac{2}{x^2+1} dx = \newline
\ln |x-1| - \ln |x+1| - 2 \arctan x + C"


2) Solution:

"\int \frac{8}{x(x^2+2)^2} dx = 
\int \frac{8x}{x^2(x^2+2)^2} dx = 
\int \frac{4}{x^2(x^2+2)^2} d(x^2)"

Using substitution "s=x^2" and denoting s to x for convenience we will receive: "\int \frac{4}{x(x+2)^2} dx"

"\frac{4}{x(x+2)^2} = \frac{A}{x} + 
\frac{B}{x+2} + \frac{C}{(x+2)^2} =
\frac{A(x+2)^2 + Bx(x+2) + Cx}{x(x+2)^2} =
\frac{x^2(A + B) + x(4A + 2B + C) + 4A}{x(x+2)^2}"

from nominator

"A+B=0 \newline
4A + 2B + C=0 \newline
4A=8"

solution of this system is: A=2 B=-2 C=-4, so

"\frac{4}{x(x+2)^2} = \frac{2}{x} - 
\frac{2}{x+2} - \frac{4}{(x+2)^2}"

and

"\int \frac{4}{x(x+2)^2} dx =
\int \frac{2}{x} dx - \int \frac{2}{x+2} dx - \int \frac{4}{(x+2)^2} dx = 
2 \ln |x| - 2 \ln |x+2| + \frac{4}{x+2} + C"